OpenStudy (anonymous):

How much is 2/3 cup plus 2/3 cup? I think it is 1 1/3 is it right?

OpenStudy (okagurao):

yeqh

OpenStudy (okagurao):

1 and one third

OpenStudy (anonymous):

Thanks

OpenStudy (superflytyguy):

4/ 3 = 1 1/3

OpenStudy (anonymous):

study for a test a few practice problems

OpenStudy (anonymous):

yeah 1 1/3

OpenStudy (anonymous):

Yes that's correct... 2/3 + 2/3 = 4/3 which equals 1 1/3

OpenStudy (anonymous):

Okay thanks I am taking the question down I will have another on so answer it